OpenStudy (anonymous):

Let f be a differentiable function such that f(-4)=12, f(9)=-4, f'(4)=-6, and f'(9)=3. The function g is differentiable and g(x)=finverse(x). What is the value of g'(-4)?

OpenStudy (anonymous):

g(x) = f^-1(x) ?

OpenStudy (anonymous):

Yes! that's what I meant

OpenStudy (anonymous):

g'(4) = 1/f' (f^-1(-4)) = 1/f'( g(-4) )

OpenStudy (anonymous):

so you need to know what g(-4) is, say c. I.e g(-4) = c and then you do 1/f'(c)

OpenStudy (anonymous):

you can find c based on the givens: f(-4)=12, f(9)=-4, f'(4)=-6, and f'(9)=3

OpenStudy (anonymous):

you get it?

OpenStudy (anonymous):

I follow that you're doing the derivative of an inverse formula where you have g'(x) =[ f-1(x) ], but when you have g'(4) = 1/f' (f^-1(-4)) = 1/f'( g(-4) ) I don't get why you put (f^-1(-4) in the denominator's argument for f'(x), or why that became the same thing as g(4) in the next step

OpenStudy (anonymous):

sorry that was a typo, it should have been g'(-4) = 1/f' (f^-1(-4)) = 1/f'( g(-4) )

OpenStudy (anonymous):

the formula for derivative of inverse function is: \[\frac{d}{dx} f^{-1} (x) = \frac{1}{f'(f^{-1} (x))}\]

OpenStudy (anonymous):

I let g(x) = f^-1(x)

OpenStudy (anonymous):

so if you replace g(x) with f^-1(x), you have: \[\frac{d}{dx} g(x) = \frac{1}{f'(g(x))}\]

OpenStudy (anonymous):

and x = -4, hence \[\frac{d}{dx} g(4) = \frac{1}{f'(g(4))}\]

OpenStudy (anonymous):

urg.. i meant -4 -.-

OpenStudy (anonymous):

so you need to know what g(-4) is. Like I said earlier, let g(-4) = c, find c, then do 1/f'(c)

OpenStudy (anonymous):

here is a hint: since g and f are inverse of each other, (-4,c) is g means (c,-4) is on f. In other words f(c) = -4

OpenStudy (anonymous):

I got g(-4) = 9, and then f'(9)=3, so doing 1/f'(c), the final answer would be 1/3? Is that right?

OpenStudy (anonymous):

so f(9) = -4, hence c = 9 1/f'(9) = 1/3. Yes, 1/3 is correct

OpenStudy (anonymous):

okay thanks! but why is the formula 1/f'(f-1x) in the first place? how did you get that?

OpenStudy (anonymous):

it's just a formula that you have to memorize to be able to do this problem